Question.... find the value of n-1 ∑ r=0 n C r / n Cr+ n C r+1 =

Question.... 
find the value of 
n-1r=0 nCr/nCr+nCr+1=

Grade:11

1 Answers

Harshit Singh
3 years ago
Dear Student

nCr+nCr+1 = n+1Cr+1
n+1Cr+1 /nCr =
n+1 / r+1
so , Our answer = r+1/n+1

Thanks